OpenStudy (anonymous):

solve the following: -11/6x = -33

OpenStudy (jlastino):

Multiply both sides of the equation by 6x 6x(-11/6x) = -33(6x) -11 = -198x Then divide by -198 to find x

OpenStudy (anonymous):

x = 18 is the answer

OpenStudy (jlastino):

No...you divided both sides of the equation by -198 so it would look like this